OpenStudy (anonymous):

Simplify the expression. 2 • 3 • 8 A. 13 B. 14 C. 48 D. 238

OpenStudy (anonymous):

Well whats 2*3?

OpenStudy (alexandervonhumboldt2):

2*3*8=6*8=???

OpenStudy (anonymous):

then times it by 8

OpenStudy (anonymous):

48

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

No problem :D

OpenStudy (alexandervonhumboldt2):

yeah